# Effect Size Estimation for Duration Recommendation in Online Experiments: Leveraging Hierarchical Models and Objective Utility Approaches

### v1

@@ 6.2 Meta-Analysis with Real Experiments
More simulations. Due to confidentiality constraints, the above dataset cannot be shared. To ensure reproducibility, we conducted a simulation study with 3000 simulated experiment trajectories at weeks 1, …4: (1) Simulate weekly sample size following beta-geometric distribution Richardson et al. (2022). The beta distribution parameters, $\alpha$ and $\beta$, are drawn from uniform distributions: $\alpha\sim Uniform(0.1,1)$, $\beta\sim Uniform(4,60)$. Each arm assumes a total of 10K customers. (2) Sample the observed effect size $d_{i,t}$ for experiment $i$ at week $t$ from the three-layer GMM model (4) with three Gaussian components $(\mu_{1},\mu_{2},\mu_{3})=(-1,0,1)$, $(\tau_{1},\tau_{2},\tau_{3})=(0.3,0.5,0.3)$, $(\pi_{1},\pi_{2},\pi_{3})=(0.2,0.6,0.2)$. Given $\sigma^{2}_{g}=500$, $\sigma_{i}^{2}$ are computed using sample size and Equation (1). (3) Total weekly opportunity cost is $c=4\times 10^{6}$, decomposed to each experiment according to their final weeks’ sample sizes. The observed effect size on last week $\{d_{i,4}\}_{i=1}^{m}$ are used to fit MLE and GMM methods. Utility-maximization uses grid-search to find the optimal effect size within the set {0.1, 0.2, …, 5}. Results in Table 4 demonstrate that the Three-layer GMM yields the most precise estimation (ground truth is 1), whereas the utility-based approach achieves the highest utility.
7 Conclusion
@@
For the GMM-based method, exploring other hierarchical models and leveraging experiment-specific features to recommend personalized effect size are important next steps. Bayesian non-parametric models (Orbanz and Teh, 2010) or structures that have been explored in bandits (Wan et al., 2021) are good starting points. For the utility-based method, we can also easily extend it to provide personalized effect size (and hence personalized duration) recommendation, by replacing the end-of-horizon posteriors of the $m$ experiments with the posterior for the target experiment at the duration recommendation time point. Besides, exploring the performance with other ATE estimators such as covariate adjusted estimators (Masoero et al., 2023) is also an interesting next step.
Appendix A Power Calculation
A.1 Frequentist Power
For two-sided tests:
$power(\delta,\sigma,\alpha)$ $=$ $1-Prob\bigg{(}t_{\alpha/2}\leq\frac{\bar{Y}_{T}-\bar{Y}_{C}}{\sigma}\leq t_{1-\alpha/2}\bigg{|}\mu_{T}-\mu_{C}=\delta\bigg{)}$
$=$ $1-Prob\bigg{(}t_{\alpha/2}-\frac{\mu_{T}-\mu_{C}}{\sigma}\leq\frac{\bar{Y}_{T}-\bar{Y}_{C}-(\mu_{T}-\mu_{C})}{\sigma}\leq t_{1-\alpha/2}-\frac{\mu_{T}-\mu_{C}}{\sigma}\bigg{|}\mu_{T}-\mu_{C}=\delta\bigg{)}$
$=$ $1-Prob\bigg{(}t_{\alpha/2}-\frac{\delta}{\sigma}\leq T\leq t_{1-\alpha/2}-\frac{\delta}{\sigma}\bigg{)}$
where $T$ is the t-statistics $\frac{\bar{Y}_{T,.}-\bar{Y}_{C,.}-(\mu_{T}-\mu_{C})}{\sigma}$. When $n\rightarrow\infty$ and variance is finite:
$power(\delta,\sigma,\alpha)$$\approx$ $1-(\Phi(z_{1-\alpha/2}-\frac{\delta}{\sigma})-\Phi(z_{\alpha/2}-\frac{\delta}{\sigma}))$
$=$ $\Phi(z_{\alpha/2}-\frac{\delta}{\sigma})+\Phi(z_{\alpha/2}+\frac{\delta}{\sigma})$
Similarly, for one-sided test:
$H_{0}\mathrel{\mathop{\ordinarycolon}}\mu\leq 0\quad H_{1}\mathrel{\mathop{\ordinarycolon}}\mu>0$
When $n\rightarrow\infty$:
$power(\delta,\sigma,\alpha)\approx\Phi(z_{\alpha}+\frac{\delta}{\sigma})$
Appendix B Methodologies
B.1 Marginal Distribution
[…]
$=$ $\frac{1}{2\pi\tau_{j}\sigma_{i}}exp\bigg{(}-\frac{1}{2}\frac{(d_{i}-\mu_{j})^{2}}{\tau_{j}^{2}+\sigma_{i}^{2}}\bigg{)}\sqrt{2\pi\frac{\tau_{j}^{2}\sigma_{i}^{2}}{\tau_{j}^{2}+\sigma_{i}^{2}}}$
$=$ $\frac{1}{\sqrt{2\pi(\tau_{j}^{2}+\sigma_{i}^{2}})}exp\bigg{(}-\frac{1}{2}\frac{(d_{i}-\mu_{j})^{2}}{\tau_{j}^{2}+\sigma_{i}^{2}}\bigg{)}$
$d_{i}|z_{i}=j\sim\mathcal{N}(\mu_{j},\tau_{j}^{2}+\sigma_{i}^{2})$
B.2 EM Algorithm
Assume $\{\sigma_{i}^{2}\}$ are known. $\bm{\theta}=\{\mu_{1},\ldots,\mu_{K},\tau_{1},\ldots,\tau_{K}\}$ and $\bm{\pi}=(\pi_{1},\ldots,\pi_{K})$ contain parameters of the each Gaussian distributions components and their corresponding weights.
We first derive a few formulas that are essential for developing the EM algorithm. Since $z$ comes from a categorical distribution:
$g(z|\bm{\pi})$ $=$ $\prod_{k=1}^{K}\pi_{k}^{\mathbb{I}[z=k]}$
The marginal distribution is as follows:
$f(d_{i}|\bm{\theta},\bm{\pi})$ $=$ $\sum_{k=1}^{K}\pi_{k}f_{k}(d_{i}|\bm{\theta})$
where $f_{k}(d_{i}|\bm{\theta})$ is the density of Gaussian distribution $\mathcal{N}(\mu_{k},\tau_{k}^{2}+\sigma_{i}^{2})$.
The conditional probability function for $z_{i}$ given the data $d_{i}$ is:
$k_{i}(z_{i}=j|d_{i},\bm{\theta},\bm{\pi})$ $=$ $\frac{\pi_{j}f_{j}(d_{i}|\bm{\theta})}{\sum_{k=1}^{K}\pi_{k}f_{k}(d_{i}|\bm{\theta})},$
Due to the independence among experiments, the joint distribution is as follows:
$p(\bm{d},\bm{z}|\bm{\theta},\bm{\pi})$ $=$ $\prod_{i=1}^{m}p_{i}(d_{i},z_{i}|\bm{\theta},\bm{\pi})$
$=$ $\prod_{i=1}^{m}f(d_{i}|\bm{\theta},\bm{\pi})k_{i}(z_{i}|d_{i},\bm{\theta},\bm{\pi})$
$=$ $\prod_{i=1}^{m}\prod_{k=1}^{K}\bigg{(}\pi_{k}f_{k}(d_{i}|\bm{\theta})\bigg{)}^{\mathbb{I}[z_{i}=k]}$
E-step. The E-step calculates the conditionally expected log-likelihood, where the expectation is conditioned on the parameters from the previous iteration step $p$ $(\bm{\theta}^{(p)},\bm{\pi}^{(p)})$ and the expectation is taken over the latent assignments.
$Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})$ $=$ $\sum_{i=1}^{m}\mathbb{E}_{z|d}\bigg{[}log\bigg{(}p_{i}(d_{i},z_{i}|\bm{\theta},\bm{\pi})\bigg{)}|\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}$
$=$ $\sum_{i=1}^{m}\sum_{k=1}^{K}\mathbb{E}_{z|d}\bigg{[}{\mathbb{I}[z_{i}=k]}|\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}\times\bigg{[}log(\pi_{k})+log(f_{k}(d_{i}|\bm{\theta}))\bigg{]}$
$=$ $\sum_{i=1}^{m}\sum_{k=1}^{K}\bigg{(}\frac{\pi_{k}^{(p)}f_{k}(d_{i}|\bm{\theta}^{(p)})}{\sum_{j=1}^{K}\ \pi_{j}^{(p)}f_{j}(d_{i}|\bm{\theta}^{(p)})}\times[log(\pi_{k})+log(f_{k}(d_{i}|\bm{\theta}))]\bigg{)}$
Line 2 to 3 above is due to $\mathbb{E}_{z|d}\bigg{[}{\mathbb{I}[z_{i}=k]}|\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}=P\bigg{[}{\mathbb{I}[z_{i}=k]}|d_{i},\bm{\theta}^{(p)},\bm{\pi}^{(p)}\bigg{]}=\frac{\pi_{k}^{(p)}f_{k}(d_{i}|\bm{\theta}^{(p)})}{\sum_{j=1}^{K}\pi_{j}^{(p)}f_{j}(d_{i}|\bm{\theta}^{(p)})}$, and we denote this quantity as $\omega_{i,k}$.
$\omega_{i,k}^{(p)}=\frac{\pi_{k}^{(p)}f_{k}(d_{i}|\bm{\theta}^{(p)})}{\sum_{j=1}^{K}\pi_{j}^{(p)}f_{j}(d_{i}|\bm{\theta}^{(p)})}$ (18)
Rewrite $Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})$ as follows:
$Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})$ $=$ $\sum_{i=1}^{m}\sum_{k=1}^{K}\bigg{(}w^{(p)}_{i,k}[log(\pi_{k})+log(f_{k}(d_{i}|\bm{\theta}))]\bigg{)}$
$=$ $\sum_{k=1}^{K}\sum_{i=1}^{m}w^{(p)}_{i,k}log(\pi_{k})+\sum_{k=1}^{K}\sum_{i=1}^{m}w^{(p)}_{i,k}\times\bigg{(}-\frac{1}{2}log(2\pi)-\frac{1}{2}log(\sigma_{i}^{2}+\tau_{k}^{2})-\frac{1}{2}\frac{(d_{i}-\mu_{k})^{2}}{\sigma^{2}_{i}+\tau_{k}^{2}}\bigg{)}$
M-step involves finding the optimal values for $\bm{\theta},\bm{\pi}$ by maximizing the log-likelihood derived in the E-step. First, we maximize $\bm{\pi}$ in $Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})$ subject to the constraint $\sum_{k=1}^{K}\pi_{k}=1$ using Lagrange Multiplier Method:
$\frac{\partial}{\partial\pi_{j}}\bigg{\{}Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})-\lambda(\sum_{k=1}^{K}\pi_{k}-1)\bigg{\}}$ $=$ $\sum_{i=1}^{m}\bigg{(}\frac{1}{\pi_{j}}\frac{\pi_{j}^{(p)}f_{j}(d_{i}|\bm{\theta}^{(p)})}{\sum_{k=1}^{K}\pi_{k}^{(p)}f_{k}(d_{i}|\bm{\theta}^{(p)})}\bigg{)}-\lambda$
Thus
$\hat{\pi}_{j}=\frac{1}{\lambda}\sum_{i=1}^{m}\bigg{(}\frac{\pi_{j}^{(p)}f_{j}(d_{i}|\bm{\theta}^{(p)})}{\sum_{k=1}^{K}\pi_{k}^{(p)}f_{k}(d_{i}|\bm{\theta}^{(p)})}\bigg{)}=\frac{1}{m}\sum_{i=1}^{m}\omega_{i,j}^{(p)}.$ (19)
Because the sum of all $\pi_{j}=1$, it follows that $\lambda=m$.
For maximizing $\bm{\theta}$ in $Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})$:
$\frac{\partial}{\partial\theta_{j}}\bigg{\{}Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})\bigg{\}}$ $=$ $\sum_{i=1}^{m}\bigg{(}\omega_{i,j}^{(p)}\frac{\frac{\partial}{\partial\theta_{j}}f_{j}(d_{i}|\theta_{j})}{f_{j}(d_{i}|\theta_{j})}\bigg{)}$
Take the derivative with respect to $\mu$ and $\tau^{2}$, and then set the derivatives equal to zero. This yields:
$\mu_{j}=\frac{\sum_{i=1}^{m}\frac{\omega_{i,j}^{(p)}d_{i}}{(\sigma^{2}_{i}+\tau^{2}_{j})}}{\sum_{i=1}^{m}\frac{\omega_{i,j}^{(p)}}{(\sigma^{2}_{i}+\tau^{2}_{j})}}$ (20)
and:
$\sum_{i=1}^{m}\omega_{i,j}^{(p)}\frac{1}{\sigma_{i}^{2}+\tau^{2}_{j}}=\sum_{i=1}^{m}\omega_{i,j}^{(p)}\frac{(d_{i}-\mu_{j})^{2}}{(\sigma_{i}^{2}+\tau^{2}_{j})^{2}}$ (21)
In summary, the EM algorithm works as the following:
E-step. From Eq 18 $\omega_{i,j}^{(p)}=\frac{\pi^{(p)}_{j}f_{j}(d_{i}|\bm{\theta^{(p)}})}{\sum_{k=1}^{K}\pi^{(p)}_{k}f_{k}(d_{i}|\bm{\theta^{(p)}})}$, where $\pi^{(p)}_{j}$ and $\bm{\theta^{(p)}}$ are all parameters from the $p$-th iteration.
M-step. Update the parameters as
$\pi_{j}$ $=$ $\frac{1}{m}\sum_{i=1}^{m}\omega_{i,j}^{(p)}$
[…]
$\mu_{j}$ $=$ $\frac{\sum_{i=1}^{m}\frac{\omega_{i,j}^{(p)}d_{i}}{(\sigma^{2}_{i}+\tau_{j}^{2})}}{\sum_{i=1}^{m}\frac{\omega_{i,j}^{(p)}}{(\sigma^{2}_{i}+\tau_{j}^{2})}}$
$\sum_{i=1}^{m}\omega_{i,j}^{(p)}\frac{1}{\sigma_{i}^{2}+\tau_{j}^{2}}$ $=$ $\sum_{i=1}^{m}\omega_{i,j}^{(p)}\frac{(d_{i}-\mu_{j})^{2}}{(\sigma_{i}^{2}+\tau_{j}^{2})^{2}}$
B.3 Penalized Three-layer GMM
One can simultaneously solve for each pair of $\mu_{j}$ and $\tau_{j}$ from $Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})$ by substituting Equation (20) to solve the value of $\tau_{j}$ first:
$Q(\bm{\theta},\bm{\pi}|\bm{\theta}^{(p)},\bm{\pi}^{(p)})+log(p_{m}(\bm{\tau^{2}}))$
$=$ $-\sum_{i=1}^{m}w^{(p)}_{i,j}\bigg{(}\frac{1}{2}log(\sigma_{i}^{2}+\tau_{j}^{2})+\frac{1}{2}\frac{(d_{i}-\mu_{j})^{2}}{\sigma^{2}_{i}+\tau_{j}^{2}}\bigg{)}-\frac{1}{m}(\frac{1}{\tau_{j}^{2}}+log(\tau_{j}^{2}))$
$L(\tau_{j}^{2})$
$=$ $-\sum_{i=1}^{m}w^{(p)}_{i,j}\bigg{(}\frac{1}{2}log(\sigma_{i}^{2}+\tau_{j}^{2})+\frac{1}{2}\frac{1}{(\tau_{j}^{2}+\sigma^{2}_{i})^{2}}\times\left[\frac{\sum_{l=1}^{m}\frac{\omega_{l,j}^{(p)}(d_{i}-d_{l})}{(\sigma^{2}_{l}+\tau_{j}^{2})}}{\sum_{l=1}^{m}\frac{\omega_{l,j}^{(p)}}{(\sigma^{2}_{l}+\tau_{j}^{2})}}\right]^{2}\bigg{)}-\frac{1}{m}(\frac{1}{\tau_{j}^{2}}+log(\tau_{j}^{2}))$
where $p_{m}(\bm{\tau^{2}})$ is the penalized function. Substitute $\operatorname*{arg\;max}_{\tau_{j}^{2}}L(\tau_{j}^{2})$ with Equation (9).
The relationship between AES and its utility value in Equation (13)using real data is illustrated in Figure LABEL:fig:optimize. As indicated by Equation (2), when all other parameters remain fixed, power exhibits an inverse proportionality to AES. Consequently, duration also displays an inverse proportionality to AES. Across the range of AES from 0.1% to 1.58%, the Utility function demonstrates an initial decline followed by an upward trend. This trend is attributed to the concurrent decrease in the decision accuracy and opportunity cost. Notably, the influence of diminishing opportunity cost becomes more dominated after the AES surpasses 1.1%.
Appendix C Simulated Data Analysis
Table 4: Simulated data analysis results (Ground truth is 1).
Pooled-MLE 0.782% 0.13% 0.1% 4.0 2.8 5.88 0.0 3.08
Two-Layer GMM 1.041% 0.13% 0.27% 3.89 2.53 5.83 0.01 3.31
Three-layer GMM 1.005% 0.13% 0.17% 3.92 2.6 5.87 0.01 3.28
Utility-maximization 1.100% 0.13% 0.37% 3.82 2.41 5.8 0.01 3.41
Table 5: Simulated data analysis results II.
Pooled-MLE 0.684% 0.0% 2.07% 3.18 2.61 9.96 0.01 7.35
Two-Layer GMM 0.917% 0.0% 4.9% 2.33 1.21 9.48 0.0 8.27
Three-layer GMM 0.933% 0.0% 5.07% 2.28 1.14 9.45 0.0 8.31
Utility-maximization 1.100% 0.0% 7.33% 1.86 0.63 9.03 0.0 8.41
We simulate 3000 experiment trajectories at week 1, 2, 3 and 4, then compare the performance of proposed two models with baseline methods in Table 4. The data generation process is as follows.
  • Simulate sample size at each week. We assume the number of customers shown in each week follows the beta-geometric model. ( Richardson et al. (2022) without the censoring in the likelihood function.) The parameters $\alpha$ and $\beta$ of the beta distribution were randomly simulated from a uniform distribution, with $\alpha\sim Uniform(0.1,1)$ and $\beta\sim Uniform(4,60)$. Assume the total number of customers is 10K in each arm.
  • Directly sample non-standardized effect size $d_{i,t}$ for every experiments $i$ at week $t$. To simply the simulation process, we make the assumption that each customer only has activity during the week in which she initially triggers. In addition, the observed effect size $d_{i,t}$ remain constant for the same $i$ throughout different weeks $t$ (These assumptions do not hold true for real-world data and are also unnecessary for our framework. They have been included solely to facilitate simulations.).
    We follow the three-layer GMM defined in Equation (5) with three Gaussian components $(\mu_{1},\mu_{2},\mu_{3})=(-1,0,1)$, $(\tau_{1},\tau_{2},\tau_{3})=(0.3,0.5,0.3)$, $(\pi_{1},\pi_{2},\pi_{3})=(0.2,0.6,0.2)$. Given $\sigma^{2}_{g}=500$, $\sigma_{i,t}^{2}$ are computed using sample size obtained from step 1 and Equation 1.
  • Total weekly opportunity cost is $c=4\times 10^{6}$, decomposed to each experiment according to their final weeks’ sample sizes.
We compare the proposed estimated AES from Three-Layer GMM and utility theory-based optimal effect size (Utility-maximization) with that of two-layer GMM and Pooled-MLE. The settings for Two-layer GMM, Three-layer GMM and Pooled-MLE are the same as Section 6.1. The observed effect size on last week $\{d_{i,4}\}_{i=1}^{m}$ are used as the training data for those three methods. Utility-maximization uses grid-search to find the optimal effect size within the set {0.1, 0.2, …, 5}.
Results in Table 4 demonstrate that the Three_layer GMM yields the most precise estimation (ground truth is 1), whereas the utility-based approach achieves the highest utility. Modify the values to $\sigma_{g}^{2}=100$ and $c=7\times 10^{6}$. Results are in Table 5 and they lead to the same conclusion.

### v2

@@ Introduction
More simulations. Due to confidentiality constraints, the above dataset cannot be shared. To ensure reproducibility, we conducted a simulation study with 3000 simulated experiment trajectories at weeks 1, …4: (1) Simulate weekly sample size following beta-geometric distribution (Richardson et al. 2022). The beta distribution parameters, $\alpha$ and $\beta$, are drawn from uniform distributions: $\alpha\sim Uniform(0.1,1)$, $\beta\sim Uniform(4,60)$. Each arm assumes a total of 10K customers. (2) Sample the observed effect size $d_{i,t}$ for experiment $i$ at week $t$ from the three-layer GMM model (4) with three Gaussian components $(\mu_{1},\mu_{2},\mu_{3})=(-1,0,1)$, $(\tau_{1},\tau_{2},\tau_{3})=(0.3,0.5,0.3)$, $(\pi_{1},\pi_{2},\pi_{3})=(0.2,0.6,0.2)$. Given $\sigma^{2}_{g}=500$, $\sigma_{i}^{2}$ are computed using sample size and Equation (1). (3) Total weekly opportunity cost is $c=4\cdot 10^{6}$, decomposed to each experiment according to their final weeks’ sample sizes. The observed effect size on last week $\{d_{i,4}\}_{i=1}^{m}$ are used to fit MLE and GMM methods. Utility-maximization uses grid-search to find the optimal effect size within the set {0.1, 0.2, …, 5}. Results in Table 3 demonstrate that the Three-layer GMM yields the most precise estimation (ground truth is 1), whereas the utility-based approach achieves the highest utility.
Conclusion
@@
For the GMM-based method, exploring other hierarchical models and leveraging experiment-specific features to recommend personalized effect size are important next steps. Bayesian non-parametric models (Orbanz and Teh 2010) or structures that have been explored in bandits (Wan, Ge, and Song 2021) are good starting points. For the utility-based method, we can also easily extend it to provide personalized effect size (and hence personalized duration) recommendation, by replacing the end-of-horizon posteriors of the $m$ experiments with the posterior for the target experiment at the duration recommendation time point. Besides, exploring the performance with other ATE estimators such as covariate adjusted estimators (Masoero, Hains, and McQueen 2023) is also an interesting next step.